# Dynamic Sparsity: Challenging Common Sparsity Assumptions for Learning World Models in Robotic Reinforcement Learning Benchmarks

### v1

@@ 6 Discussion
Our study provides insights into the role of sparsity in learning the dynamics of classical reinforcement learning (RL) environments and its implications for model-based reinforcement learning (MBRL). First, we observed that globally sparse causal structures, as indicated by consistently zero Jacobian elements, are rare across most environments. While some environments, such as ReacherHard and CheetahRun, exhibited limited global sparsity, the majority showed dense interactions between state and action variables. This suggests that enforcing strong global sparsity priors in learned dynamics models may not be universally beneficial. Interestingly, we found evidence of state-dependent sparsity, where the causal structure of the dynamics changes based on the current state. For example, in the FingerSpin environment, sparsity was higher when the finger was not in contact with the object and lower during interactions.
In our learning experiments, we observed that while a two-layer MLP trained to approximate could recover some sparsity in its predictions, the induced sparsity remained far below the target sparsity. Using MAE-based Jacobian losses slightly improved next-state prediction performance compared to MSE-based losses, though the differences were not statistically significant. Including ground-truth Jacobian losses increased induced sparsity but did not fully capture the sparsity present in the ground-truth dynamics, leaving the network’s predictions more entangled.
@@ 7 Conclusion
While sparsity priors hold promise for improving sample efficiency and generalization, their effectiveness depends on alignment with the true structure of the environment. Overly strong or misaligned priors could hinder learning by blocking important interactions. Our findings indicate that sparsity is indeed present in many reinforcement learning environments, but requires modelling in a state dependent way. Additionally, we observed that changes in sparsity often occur in a temporally sparse manner, with periods of stable sparsity interspersed with abrupt transitions such as contact events or phase changes in locomotion. Our results also indicate that naive MLP implementations are insufficient to fully capture and exploit these sparsity structures. These findings highlight the need for further development of dynamics model architectures that can explicitly model sparsity and thus dynamically adapt their causal structure based on state and time, thereby improve generalization and interpretability.
@@
Appendix A1 Selection of Agents
We generate a dataset of trajectories per environment using pre-trained PPO agents (Raffin et al. 2021). This is necessary due to publicly available datasets not containing the ground-truth Jacobians. In total we collect million steps per environment, using the learned stochastic agent policy. The mean reward achieved across the full episodes in our dataset is shown in Table˜A1.1.
@@
A2.1 Hyperparameter Search
@@
We perform the same procedure a final time to find a choice of activation function between the Exponential Linear Units (ELU) (Clevert, Unterthiner, and Hochreiter 2015) and the Rectified Linear Unit (ReLU) (Fukushima 1969). The resulting ranks and choice are shown in Table˜A2.13.
@@
A3.1 Cumulative Sparsity Heatmaps

### v2

@@ 6 Discussion
Our study provides insights into the role of sparsity in learning the dynamics of classical reinforcement learning (RL) environments and its implications for model-based reinforcement learning (MBRL). First, we observed that globally sparse causal structures, as indicated by consistently zero Jacobian elements, are rare across most environments: While some environments, such as ReacherHard and CheetahRun, exhibited limited global sparsity, the majority showed dense interactions between state and action variables. This suggests that enforcing strong global sparsity priors in learned dynamics models may not be universally beneficial. Interestingly, we found evidence of state-dependent sparsity i.e. the causal structure of the dynamics changes based on the current state. For example, in the FingerSpin environment, sparsity was higher when the finger was not in contact with the object and lower during interactions.
In our learning experiments, we observed that while a two-layer MLP trained to approximate could recover some sparsity in its predictions, the induced sparsity remained far below the target sparsity. Using MAE-based Jacobian losses slightly improved next-state prediction performance compared to MSE-based losses, though the differences were not statistically significant. Including ground-truth Jacobian losses increased induced sparsity but did not fully capture the sparsity present in the ground-truth dynamics, leaving the network’s predictions more entangled. This likely leads to reduced generalization and more compounding errors during multistep prediction.
@@ 7 Conclusion
While sparsity priors hold promise for improving sample efficiency and generalization, their effectiveness depends on alignment with the true structure of the environment. Overly strong or misaligned priors could hinder learning by blocking important interactions. Our findings indicate that sparsity is indeed present in many reinforcement learning environments, but requires modelling in a state dependent way. Additionally, we observed that changes in sparsity often occur in a temporally sparse manner, with periods of stable sparsity interspersed with abrupt transitions such as contact events or phase changes in locomotion. Our results also indicate that naive MLP implementations are insufficient to fully capture and exploit these sparsity structures. These findings highlight the need for further development of dynamics model architectures that can explicitly model sparsity and thus dynamically adapt their causal structure based on state and time, thereby improving generalization and interpretability.
Acknowledgements
We thank Christof Beck, Ananth Rachakonda, and Henri Geiß for their helpful comments on earlier drafts.
This research was partially funded by the Austrian Science Fund (FWF):
I 5755-N (ELSA), and by
the Autonomous Province of Bolzano-Bozen - South
Tyrol under Funding Agreement 10/2024, Abstractron.
@@ Acknowledgements
Appendix A1 Agents & Data Collection
We generate a dataset of trajectories per environment using pre-trained PPO agents based on the implementation by (Raffin et al. 2021), more specifically following the colored noise methodology proposed by (Hollenstein, Martius, and Piater 2024). Multiple agents were trained using the hyperparameter configurations described in Table˜A1.2. For each environment, the best performing agent was selected. The mean reward achieved across the full episodes in our dataset is shown in Table˜A1.2.
Dataset generation then uses the selected best agents. Dataset generation is necessary because there are no publicly available datasets containing ground-truth Jacobians of the transition functions. For the dataset, we collect a total of million steps per environment, using the learned stochastic agent policy (i.e., sampling actions from the parameterized-Gaussian policy). To improve fidelty MJX and JAX where switched to float64 mode. The MLP parameters in the learning experiments were represented in float32.
| Param. | Value | Default |
| lr | 5e-05 | 0.0003 |
| n-steps | 64 | 2048 |
| batch-size | 64 | |
| n-epochs | 10 | |
| gamma | 0.99 | |
| gae-lambda | [0.0, 0.95] | 0.95 |
| clip-range | 0.2 | |
| normalize-advantage | True | |
| ent-coef | 0 | |
| vf-coef | 0.5 | |
| use-sde | False | |
| sde-sample-freq | -1 | |
| max-grad-norm | 0.5 | |
| stats-window-size | 100 | |
| eval-episodes | 50 | |
| eval-freq | 10240 | |
| n-envs | 128 | |
| total-timesteps | 2.04801e+06 | |
| seed | [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] | |
| noise-color | [0.0, 0.5, 1.0, 2.0] | |
| n-eval-envs | 10 |
@@
A2.1 Hyperparameter Search for Dynamics Models
@@
We perform the same procedure a final time to find a choice of activation function between the Exponential Linear Units (ELU) (Clevert, Unterthiner, and Hochreiter 2016) and the Rectified Linear Unit (ReLU) (Fukushima 1969). The resulting ranks and choice are shown in Table˜A2.13.
@@
A3.1 Sparsity Heatmaps Showing Zero-Duration Proportions
Sparsity heatmaps showing the duration for which the elements in the state and action Jacobians were zero as a proportion of the total environment steps in a rollout. The following heatmaps are for environments which were not included in the paper for both state and action Jacobians in the Figure˜A3.1 and Figure˜A3.2